Formatting and Submission Instructions for the Nineteenth International Conference on Machine Learning

Pat Langley langley@isle.org Institute for the Study of Learning and Expertise, 2164 Staunton Court, Palo Alto, CA 94306 USA

Claude-Nicolas Fiechter fiechter@rtna.daimlerchrysler.com Mehmet Göker goker@rtna.daimlerchrysler.com DaimlerChrysler Research and Technology Center, 1510 Page Mill Road, Palo Alto, CA 94304 USA

Cynthia Thompson cthomp@csli.stanford.edu Computational Learning Laboratory, Center for the Study of Language and Information, Stanford University, Stanford, CA 94305 USA

Andrea Danyluk andrea@cs.williams.edu Department of Computer Science, Williams College, Williamstown, MA 01267 USA

Claude Sammut claude@cse.unsw.edu.au School of Computer Science and Engineering, University of New South Wales, Sydney, Australia

Abstract

The paper abstract should begin in the left column, 0.4 inches below the final address. The heading ‘Abstract’ should be centered, bold, and in 11 point type. The abstract body should use 10 point type, with a vertical spacing of 11 points, and should be indented 0.25 inches more than normal on left-hand and right-hand margins. Insert 0.4 inches of blank space after the body. Keep your abstract brief, limiting it to one paragraph and no more than six or seven sentences.

1. Format of the Paper

All submissions should follow the same format to ensure the printer can reproduce them without problems and to let readers more easily find the information that they desire.

1.1 Length and Dimensions

Papers must not exceed eight (8) pages, including all figures, tables, references, and appendices. We will return to the authors any submissions that exceed this page limit or that diverge significantly from the format specified herein.

The text of the paper should be formatted in two columns, with an overall width of 6.75 inches, length of 9.0 inches, and 0.25 inches between the columns. The left margin should be 0.75 inches and the top margin 1.0 inch (2.54 cm). The right and bottom margins will depend on whether you print on US letter or A4 paper.

The paper body should be set in 10 point type with a vertical spacing of 11 points. Please use Times Roman typeface throughout the text.

1.2 Title and Author Information

The paper title should be set in 14 point bold type and centered between two horizontal rules that are 1 point thick, with 1.0 inch between the top rule and the top edge of the page. Capitalize the first letter of content words and put the rest of the title in lower case.

Author information should start 0.3 inches below the bottom rule surrounding the title. The authors’ names should appear in 10 point bold type, electronic mail addresses in 10 point small capitals, and physical addresses in ordinary 10 point type.

Each author’s name should be flush left, whereas the email address should be flush right on the same line. The author’s physical address should appear flush left on the ensuing line, on a single line if possible. If successive authors have the same affiliation, then give their physical address only once.

1.3 Partitioning the Text

You should organize your paper into sections and paragraphs to help readers place a structure on the material and understand its contributions.

1.3.1 Sections and Subsections

Section headings should be numbered, flush left, and set in 11 pt bold type with the content words capitalized. Leave 0.25 inches of space before the heading and 0.15 inches after the heading.

Similarly, subsection headings should be numbered, flush left, and set in 10 pt bold type with the content words capitalized. Leave 0.2 inches of space before the heading and 0.13 inches afterward.

Finally, subsubsection headings should be numbered, flush left, and set in 10 pt small caps with the content words capitalized. Leave 0.18 inches of space before the heading and 0.1 inches after the heading. Please use no more than three levels of headings.

1.3.2 Paragraphs and Footnotes

Within each section or subsection, you should further partition the paper into paragraphs. Do not indent the first line of a given paragraph, but insert a blank line between succeeding ones.

You can use footnotes^{1} to provide readers with additional information about a topic without interrupting the flow of the paper. Indicate footnotes with a number in the text where the point is most relevant. Place the footnote in 9 point type at the bottom of the column in which it appears. Precede the first footnote in a column with a horizontal rule of 0.8 inches.^{2}

1.4 Figures

You may want to include figures in the paper to help readers visualize your approach and your results. Such artwork should be centered, legible, and separated from the text. Lines should be dark and at least 0.5 points thick for purposes of reproduction, and text should not appear on a gray background.

Label all distinct components of each figure. If the figure takes the form of a graph, then give a name for each axis and include a legend that briefly describes each curve. However, do not include a title above the figure, as the caption already serves this function.

Figure 1. Steps in the computational discovery process at which the developer can influence system behavior.

Table 1. Classification accuracies for naive Bayes and flexible Bayes on various data sets.

Data set  Naive  Flexible Better?
Breast  95.9± 0.2 96.7± 0.2  √
Cleveland 83.3± 0.6 80.0± 0.6  ×
Credit  74.8± 0.5 78.3± 0.6
Glass2  61.9± 1.4 83.8± 0.7  √
Horse  73.3± 0.9 69.7± 1.0  ×
Meta  67.1± 0.6 76.5± 0.5  √
Pima  75.1± 0.6 73.9± 0.5
Vehicle  44.9± 0.6 61.5± 0.4  √

Number figures sequentially, placing the figure number and caption after the graphics, with at least 0.1 inches of space before the caption and 0.1 inches after it, as in Figure 1. The figure caption should be set in 9 point type and centered unless it runs two or more lines, in which case it should be flush left. You may float figures to the top or bottom of a column, and you may set wide figures across both columns, but always place two-column figures at the top or bottom of the page.

1.5 Tables

You may also want to include tables that summarize material. Like figures, these should be centered, legible, and numbered consecutively. However, place the title above the table with at least 0.1 inches of space before the title and the same after it, as in Table 1. The table title should be set in 9 point type and centered unless it runs two or more lines, in which case it should be flush left.

Tables contain textual material that can be typeset, as contrasted with figures, which contain graphical material that must be drawn. Specify the contents of each row and column in the table’s topmost row. Again, you may float tables to a column’s top or bottom, and set wide tables across both columns, but place twocolumn tables at the top or bottom of the page.

1For the sake of readability, footnotes should be complete sentences.

2Multiple footnotes can appear in each column, in the same order as they appear in the text, but spread them across columns and pages if possible.

1.6 Citations and References

Please use APA reference format regardless of your formatter or word processor. If you rely on the L^{A}TEX bibliographic facility, use mlapa.sty and mlapa.bst at the ICML-2002 web site to obtain this format.

Citations within the text should include the authors’ last names and year. If the authors’ names are included in the sentence, place only the year in parentheses, as in Jones and VanLehn (1994), but otherwise place the entire reference in parentheses with the authors and year separated by a comma (Jones & Van-Lehn, 1984).

List multiple references alphabetically and separate them by semicolons (Jones & VanLehn, 1984; Veloso & Carbonell, 1993). Use the ‘et al.’ construct only for citations with four or more authors or after listing all authors to a publication in an earlier reference.

Use an unnumbered first-level section heading for the references, and use a hanging indent style, with the first line of the reference flush against the left margin and subsequent lines indented by 10 points. The references at the end of this document give examples for journal articles, conference publications, book chapters, books, edited volumes, technical reports, and dissertations.

Alphabetize references by the surnames of the first authors, with single author entries preceding multiple author entries. Order references for the same authors by year of publication, with the earliest first.

2. Electronic Submission

ICML-2002 will rely heavily on electronic formats for submission and review. We assume that nearly all authors will have access to standard software for word processing, electronic mail, and ftp file transfer. Authors who do not have such access should send email

with their concerns to icml2002@cse.unsw.edu.au.

2.1 Templates for Papers

Electronic templates for producing the camera-ready copy are available for several major word processors, including L^{A}TEX and Microsoft Word. Templates are accessible on the World Wide Web at:

 http://www.cse.unsw.edu.au/∼icml2002/format.html

Send questions about these electronic templates to

                                                icml2002@cse.unsw.edu.au.

2.2 Initial Submission of Papers

Submission to ICML-2002 will be entirely electronic. To submit a paper, please go to the submission web site:

                                                    http://www.cse.unsw.edu.au/∼icml2002/cyberchair.html

The deadline for submissions to ICML-2002 is Friday, 1 February 2002. If your submission does not reach us by this date, it will not be considered for publication.

To ensure our ability to print submissions, authors must provide their manuscripts in postscript or pdf format. If you are preparing your paper in Word, please use the Apple LaserWriter 16/600 PS driver to ensure its printability in other environments. If you cannot deliver a postscript or pdf file electronically due to exceptional conditions, send email

                                                to icml2002@cse.unsw.edu.au to discuss alternative

means of delivery.

ICML-2002 allows simultaneous submission to other conferences, provided this fact is clearly indicated on the submission form. Accepted papers will appear in the conference proceedings only if withdrawn from other conferences. Simultaneous submissions that are not clearly specified as such will be rejected.

2.3 Submitting Revised Papers

Final versions of papers accepted for publication, and resubmissions of conditionally accepted papers, should follow the same format and naming convention as initial submissions.

The deadline for revised copies of conditionally accepted papers is Monday, 15 April 2002. To submit a revised paper for final review, please go to:

                                                        http://www.cse.unsw.edu.au/∼icml2002/submit/

Camera-ready copy of all accepted papers (conditional and unconditional) must be received by the publisher by Friday, 19 April 2002. Please see the Author Kit that has been made available from the ICML-2002 homepage:

                                                            http://www.cse.unsw.edu.au/∼icml2002/

Please note that the instructions include a Permission to Publish form that must be signed and sent to the publisher.

After submitting your final copy to the publisher, please be sure to submit a copy to the conference submission site as well:

http://www.cse.unsw.edu.au/∼icml2002/submit/

If your revised paper does not reach us by the due date, it will not be included in the proceedings.

Acknowledgements

Please place your acknowledgements in an unnumbered section at the end of the paper. Typically, this will include thanks to reviewers who gave useful comments, to colleagues who contributed to the ideas, and to funding agencies and corporate sponsors that provided financial support. This document was modified by Andrea Danyluk from the file made available by Pat Langley for ICML-2K.

References

Aha, D. W. (1990). A study of instance-based algorithms for supervised learning tasks: Mathematical, empirical, and psychological evaluations. Doctoral dissertation, Department of Information & Computer Science, University of California, Irvine.

Fisher, D. H. (1989). Noise-tolerant conceptual clustering. Proceedings of the Eleventh International Joint Conference on Artificial Intelligence (pp. 825– 830). San Francisco: Morgan Kaufmann.

Jones, R. M., & VanLehn, K. (1994). Acquisition of children’s addition strategies: A model of impassefree, knowledge-level learning. Machine Learning, 16 , 11–36.

Langley, P. (1995). Elements of machine learning. San Francisco: Morgan Kaufmann.

Maloof, M. A., Langley, P., Binford, T. O., & Sage, S. (1998). Improving rooftop detection in aerial images through machine learning (Technical Report 98-1). Institute for the Study of Learning and Expertise, Palo Alto, CA.

Shrager, J., & Langley, P. (Eds.). (1990). Computational models of scientific discovery and theory formation. San Francisco: Morgan Kaufmann.

Veloso, M. M., & Carbonell, J. G. (1993). Toward scaling up machine learning: A case study with derivational analogy in Prodigy. In S. Minton (Ed.), Machine learning methods for planning. San Francisco: Morgan Kaufmann.